# RiemannLoRA: A Unified Riemannian Framework for Ambiguity-Free LoRA Optimization

### v1

@@ Abstract
Low-Rank Adaptation (LoRA) has become a widely adopted standard for parameter-efficient fine-tuning of large language models (LLMs), significantly reducing memory and computational demands. However, challenges remain, including finding optimal initialization strategies or mitigating overparametrization in low-rank matrix factorization. In this work, we propose a novel approach that addresses both of the challenges simultaneously within a unified framework. Our method treats a set of fixed-rank LoRA matrices as a smooth manifold. Considering adapters as elements on this manifold removes overparametrization, while determining the direction of the fastest loss decrease along the manifold provides initialization. Special care is taken to obtain numerically stable and computationally efficient implementation of our method, using best practices from numerical linear algebra and Riemannian optimization. Experimental results on LLM and diffusion model architectures demonstrate that RiemannLoRA consistently improves both convergence speed and final performance over standard LoRA and its state-of-the-art modifications.
@@ 1 Introduction
Large language models (LLMs) have demonstrated remarkable capabilities across a wide range of natural language processing tasks Brown et al. (2020); Touvron et al. (2023a, b). However, the computational and storage costs associated with training and deploying such models at scale pose significant challenges. To reduce these costs, parameter-efficient fine-tuning techniques such as low-rank adaptation (LoRA) Hu et al. (2022) have emerged as a practical solution. LoRA enables efficient adaptation of pre-trained models by embedding learnable low-rank matrices into specific weight updates, allowing most of the original parameters to remain frozen. In particular, the main idea of LoRA is to fine-tune a pretrained model using a rank-$r$ correction matrix $\Delta W$:
@@
Due to its widespread adoption and utility, numerous efforts have been made to further enhance the quality of low-rank fine-tuning. Various ambiguities have inspired multiple techniques that leverage them for performance gains. The two primary sources of ambiguity in low-rank adaptation are as follows. First, we can improve LoRA initialization by finding better starting matrices $A_{0}\in\mathbb{R}^{m\times r}$ and $B_{0}\in\mathbb{R}^{n\times r}$. This is done by replacing the original weight matrix $W$ with a modified version $W^{\prime}$ using the following splitting:
$W=\underbrace{(W-A_{0}B_{0}^{\top})}_{W^{\prime}}+A_{0}B_{0}^{\top}=W^{\prime}+A_{0}B_{0}^{\top},$ (1)
with $(A_{0},B_{0})$ chosen based on low-rank approximation of weight matrices or gradients (Meng et al., 2024; Wang et al., 2025, 2024) and one of the matrices is chosen to be zero is standard LoRA. On the other hand, there is a problem in overparametrization of low-rank matrices during training. Indeed, for any $U_{1}\in\mathbb{R}^{m\times r}$,$V_{1}\in\mathbb{R}^{n\times r}$ and any invertible matrix $S\in\mathbb{R}^{r\times r}$, we may write:
$AB^{\top}=\widetilde{A}{\widetilde{B}}^{\top},\quad\text{where}\quad\widetilde{A}=AS,\quad\widetilde{B}=BS^{-\top}.$ (2)
Widely used optimization methods such SGD (Robbins and Monro, 1951), Adam (Kingma and Ba, 2014), Adagrad (Duchi et al., 2011), RMSProp (Tieleman, 2012) neither account for nor exploit these inherent ambiguities.
In this work, we propose a novel framework that naturally eliminates all ambiguities within a unified approach. To the best of our knowledge, this is the first such attempt. In particular, we introduce a new method called Riemannian LoRA (RiemannLoRA). This approach replaces the standard Euclidean optimization of low-rank matrices with one on Riemannian manifold of fixed-rank matrices, thereby respecting the intrinsic geometry of low-rank matrices. On this manifold, every point is considered as an element of $\mathbb{R}^{m\times n}$, so it does not depend on a particular parametrization choice (2). The Riemannian framework also allows us to initialize the method in a geometrically meaningful way – by choosing the splitting that yields the fastest loss decrease along the manifold.
The contributions of this work are as follows.
  • We propose a new framework, called RiemannLoRA, which natively resolves ambiguities of previous LoRA-based approaches using Riemannian optimization. It combines two essential ingredients: initialization (Section 4.2) and parametrization-free optimization (Sections 4.14.3), that is based on inversion-free robust formulas and reliable matrix decompositions.
  • We propose an efficient way to initialize matrices by using a randomized SVD algorithm with implicit matrix multiplication (Section 4.2).
@@
  • We showcase the performance of our framework for fine-tuning LLMs and in subject-driven generation using diffusion models. Among positive effects that we observe are: boost in target metrics, improved convergence, and reduction of variance (Section 5).
@@ 2 Related Work
The problem of an optimal initial guess selection for low-rank LLM adaptation has been addressed in a sequence of works: the authors Meng et al. (2024, PiSSA) have suggested a heuristic that involves using a low-rank truncated SVD of pretrained parameters as an initial point for LoRA and its orthogonal complement as frozen layer’s parameters, so that the tuning process starts without changing the starting value of the loss function. A similar approach was implemented by Wang et al. (2025, MiLoRA) with the main difference of optimizing the smallest singular components of unadapted parameter matrix. A context-aware initialization was considered in (Yang et al., 2024, CorDA) and (Parkina and Rakhuba, 2025, COALA) proposes a numerically robust inversion-free framework for low-rank weighted approximations for this setting. Another idea is to initialize LoRA with a subset of left and right singular vectors of a doubled-rank truncated SVD of the loss function gradient at the starting parameters, proposed by Wang et al. (2024, LoRA-GA). We show direct connection of this method to our Riemannian initialization strategy and propose how to additionally significantly accelerate the computation of SVD using our approach. Attempting to overcome the asymmetry in the initialization of vanilla LoRA fine-tuning process, (Hayou et al., 2024, LoRA+) introduced a scale-free step size selection for LoRA factors.
@@
Riemannian optimization is widely used for algorithms on matrix manifolds and allows for exploiting task geometry or imposing additional constraints. For example, a Riemannian solution for the extreme eigenpairs search problem was described in Absil et al. (2009); Baker (2008), a matrix completion task, which is common in collaborative filtering for recommender systems, via optimization on the fixed-rank manifold (Vandereycken, 2013)), a Riemannian approach on the manifold of matrices with orthonormal columns (the Stiefel manifold) was used by Wisdom et al. (2016) for diminishing the problem of vanishing and exploding gradients in recurrent neural networks, etc. The book Trendafilov and Gallo (2021) also presents a comprehensive description of useful manifolds for solutions of the data science problems. For the deeper understanding of applied differential geometry techniques see the books Absil et al. (2009) and Boumal (2023).
@@
The idea of using Riemannian optimization has recently started to emerge for the large language models. For example, the fine-tuning of LLMs with the help of the Stiefel manifold was considered in the work Hu et al. (2024). The authors of (Zhang and Pilanci, 2024) introduced the Riemannian inspired modification of Adam. The authors of Mo et al. (2025, LORO) applied the Riemannian optimization techniques for pretraining LLMs on the fixed-rank manifold. Parametrization that is used in our work can potentially help in this setting as well, by additionally avoiding potential overheads and instabilities, arising due the explicit inversion of Gram matrices.
3 Riemannian optimization on a fixed-rank manifold
If we only fine-tune the parameter matrix of one layer, the optimization problem becomes:
$\mathcal{L}(W+AB^{\top})\to\min_{A,B}.$
In the Riemannian optimization approach, we treat each low-rank matrix $\Delta W=AB^{\top}$ as an element of the manifold of fixed-rank matrices $\mathcal{M}_{r}$ (formal definitions are provided later):
$\mathcal{L}(W+\Delta W)\to\min_{\Delta W\in\mathcal{M}_{r}}.$
Consequently, rather than relying on a specific parameterization, we traverse from one point on $\mathcal{M}_{r}$ to another. Thus, in exact arithmetic, all valid representations of $\Delta W\in\mathcal{M}_{r}$ lead to equivalent methods. Nevertheless, the choice of parameterization is important for computational efficiency and numerical stability, and we pay attention to these aspects as well.
3.1 Fixed-rank manifold
Let $r<\min(m,n)$ be a matrix rank, and consider the set
$\mathcal{M}_{r}=\left\{X\in\mathbb{R}^{m\times n}|\mathrm{rank}\,(X)=r\right\}\subset\mathbb{R}^{m\times n}.$ (3)
$\mathcal{M}_{r}$ forms a smooth embedded into linear space $\mathbb{R}^{m\times n}$ manifold (Lee, 2003, Example 8.14) of dimension
$\mathrm{dim}\,\mathcal{M}_{r}=(m+n)r-r^{2}.$
Every rank-$r$ matrix $X\in\mathcal{M}_{r}$ can be represented using certain full-rank matrices $A\in\mathbb{R}^{m\times r},B\in\mathbb{R}^{n\times r}$ as
$X=AB^{\top}.$
By orthogonalizing the columns of $A$ or $B$, we can parametrize the manifold of fixed-rank in the following two ways:
$X=A_{L}B^{\top}=AB^{\top}_{R},\quad\forall X\in\mathcal{M}_{r},$ (4)
where $A,A_{L}\in\mathbb{R}^{m\times r}$, $B,B_{R}\in\mathbb{R}^{n\times r}$ and $A_{L},B_{R}$ have orthonormal columns
$A_{L}^{\top}A_{L}=I_{r},\quad B_{R}^{\top}B_{R}=I_{r},\quad I_{r}\in\mathbb{R}^{r\times r}.$
This representation is a particular case of the representation of elements of the fixed tensor train rank manifold (Holtz et al., 2012) and can be achieved, for example, with the help of the QR decomposition. Several alternative parameterizations of the fixed-rank matrix manifold exist (see (Boumal, 2023, Chap. 7.5)), though we do not employ them in this work.
The smoothness of the embedded manifold ensures that, locally, it is well-approximated by a linear subspace called tangent space, which resembles local linearization of smooth functions. If $X\in\mathcal{M}_{r}$ then the tangent space at that point can be represented as
$\mathcal{T}_{X}\mathcal{M}_{r}=\left\{\xi\big{|}\ \xi=\dot{A}B_{R}^{\top}+A_{L}\dot{B}^{\top},\ X=A_{L}B^{\top},\ \dot{A}^{\top}A_{L}=\mathbf{0}_{r}\right\}\subset\mathbb{R}^{m\times n}.$ (5)
Just as points on a fixed-rank manifold can be represented as a pair of matrices $(A,B)$, tangent vectors can also be represented in a similar format, though with the ranks twice as large. In particular, for any $\xi\in\mathcal{T}_{X}\mathcal{M}_{r}$:
$\xi=\begin{bmatrix}\dot{A}&A_{L}\end{bmatrix}\begin{bmatrix}B_{R}\\ \dot{B}\end{bmatrix}^{\top},\quad\mathrm{rank}\,\xi\leq 2r.$ (6)
Interestingly, the point $X\in\mathcal{M}_{r}$ itself lies in the $\mathcal{T}_{X}\mathcal{M}_{r}$ with $\dot{A}=0,\dot{B}=B$.
@@ 3.1 Fixed-rank manifold
Given a matrix $Z\in\mathbb{R}^{m\times n}$, its orthogonal projection $P_{\mathcal{T}_{X}\mathcal{M}_{r}}\,Z$ onto the tangent space $\mathcal{T}_{X}\mathcal{M}_{r}$ (with the parameterization given in (5)) can be computed as follows:
$\begin{split}&P_{\mathcal{T}_{X}\mathcal{M}_{r}}\,Z=\dot{A}_{Z}B_{R}^{\top}+A_{L}\dot{B}_{Z}^{\top},\quad\forall Z\in\mathbb{R}^{m\times n},\\ &\dot{A}_{Z}=(I-A_{L}A_{L}^{\top})ZB_{R},\quad\dot{B}_{Z}=Z^{\top}A_{L}.\end{split}$ (7)
3.2 Riemannian optimization techniques
@@
Let $\mathcal{M}_{r}\subset\mathbb{R}^{m\times n}$ be a smooth submanifold and $F\colon\mathbb{R}^{m\times n}\to\mathbb{R}$ a differentiable function with the Euclidean gradient $\nabla F\in\mathbb{R}^{m\times n}$. Within the Riemannian optimization framework, we solve the following task via gradient-based optimization

### v2

# LoRA meets Riemannion: Muon Optimizer for Parametrization-independent Low-Rank Adapters

@@ Abstract
This work presents a novel, fully Riemannian framework for Low-Rank Adaptation (LoRA) that geometrically treats low-rank adapters by optimizing them directly on the fixed-rank manifold. This formulation eliminates the parametrization ambiguity present in standard Euclidean optimizers. Our framework integrates three key components to achieve this: (1) we derive Riemannion, a new Riemannian optimizer on the fixed-rank matrix manifold that generalizes the recently proposed Muon optimizer; (2) we develop a Riemannian gradient-informed LoRA initialization, and (3) we provide an efficient implementation without prominent overhead that uses automatic differentiation to compute arising geometric operations while adhering to best practices in numerical linear algebra. Comprehensive experimental results on both LLM and diffusion model architectures demonstrate that our approach yields consistent and noticeable improvements in convergence speed and final task performance over both standard LoRA and its state-of-the-art modifications.
@@ 1 Introduction
Large language models (LLMs) have demonstrated remarkable capabilities across a wide range of natural language processing tasks Brown et al. (2020); Touvron et al. (2023a; b). However, the computational and storage costs associated with training and deploying such models at scale pose significant challenges. To reduce these costs, parameter-efficient fine-tuning techniques such as low-rank adaptation (LoRA) Hu et al. (2022) have emerged as a practical solution. LoRA enables efficient adaptation of pre-trained models by embedding learnable low-rank matrices into specific weight updates, allowing most of the original parameters to remain frozen. In particular, the main idea of LoRA is to fine-tune a pretrained model using a rank-$r$ correction matrix $\Delta W$:
@@
Despite its efficiency, the dominant practice of optimizing the LoRA factors $(A,B)$ with Euclidean optimizers such as SGD (Robbins & Monro, 1951), Adam (Kingma & Ba, 2014), Adagrad (Duchi et al., 2011), RMSProp (Tieleman, 2012), etc. that misaligned with the geometry of the low-rank constraint. The same update $\Delta W$ can be represented by infinitely many factorizations: for any $A\in\mathbb{R}^{m\times r}$, $B\in\mathbb{R}^{n\times r}$ and any invertible matrix $S\in\mathbb{R}^{r\times r}$, we may write:
$\Delta W=AB^{\top}=\widetilde{A}{\widetilde{B}}^{\top},\quad\text{where}\quad\widetilde{A}=AS,\quad\widetilde{B}=BS^{-\top}.$ (1)
Ideally, training should be reparameterization (transformation) invariant: the update to $\Delta W$ must not depend on which factorization $(A,B)$ is used (Yen et al., 2024). Empirically, this lack of invariance manifests as unbalanced learning where one factor dominates and the other stalls, fragile hyperparameter sensitivity, and path-dependent solutions. These issues have prompted geometry-aware formulations. Riemannian treatments of low-rank models operate on the fixed-rank manifold rather than the ambient factor space, projecting gradients to the tangent space and retracting back to the manifold. Such steps can be implemented efficiently when $r\ll\min\{m,n\}$ and avoid forming full-size matrices. Within the LoRA literature, existing Riemannian approaches either use standard SGD-type optimizers Mo et al. (2025, LORO) or rely on the Adam (Zhang & Pilanci, 2024) optimizer for auxiliary matrices within the chosen parameterization, which deviate them from the Riemannian framework and introduce dependence on parameterization.
In this work, we introduce a fully Riemannian framework for training LoRA that optimizes the adapter $X\!=\!\Delta W$ directly on the fixed-rank matrix manifold
$\mathcal{M}_{r}\;=\;\{X\in\mathbb{R}^{m\times n}:\mathrm{rank}(X)=r\},$
eliminating factorization ambiguity by construction. Central to our approach is Riemannion, a new Riemannian optimizer on $\mathcal{M}_{r}$ that generalizes the recently proposed Muon optimizer (Jordan et al., 2024) to the fixed-rank setting. In contrast to prior Riemannian LoRA variants that port Adam-like mechanics to the manifold with ad hoc choices, our design inherits Muon’s geometry-aligned normalization, yielding transformation invariance of the learned update. We further propose a Riemannian gradient–informed initialization that places the initial adapter at a good location on $\mathcal{M}_{r}$, and we provide a practical, low-overhead implementation that assembles projections, retractions, and vector transports via automatic differentiation, also following best practices from numerical linear algebra. Extensive experiments on LLM and diffusion architectures show consistent gains in convergence speed and final task performance over standard LoRA and recent state-of-the-art modifications.
Our contributions are as follows:
  • Riemannion: Muon on the fixed-rank manifold. We derive Riemannion, the first optimizer that generalizes Muon to the manifold $\mathcal{M}_{r}$ of fixed rank matrices.
  • Riemannian gradient-informed initialization. We propose an initialization strategy which yields best alignment between the initial Riemannian gradient and the Euclidean gradient. We also propose an efficient way for this strategy by using a randomized SVD algorithm with implicit matrix multiplication (Section 5). Finally, we show the connection of this initialization to LoRA-GA.
  • Efficient implementation with automatic differentiation. We pay special attention to numerical implementation to make the method robust without any prominent overhead compared to vanilla LoRA at small ranks.
@@
  • Comprehensive empirical validation. We showcase the performance of our framework for fine-tuning LLMs and in subject-driven generation using diffusion models. Among positive effects that we observe are: boost in target metrics, improved convergence, and reduction of variance.
@@ 2 Related Work
The problem of an optimal initial guess selection for low-rank LLM adaptation has been addressed in a sequence of works: the authors Meng et al. (2024, PiSSA) have suggested a heuristic that involves using a low-rank truncated SVD of pretrained parameters as an initial point for LoRA and its orthogonal complement as frozen layer’s parameters, so that the tuning process starts without changing the starting value of the loss function. A similar approach was implemented by Wang et al. (2025, MiLoRA) with the main difference of optimizing the smallest singular components of unadapted parameter matrix. A context-aware initialization was considered in (Yang et al., 2024, CorDA) and (Parkina & Rakhuba, 2025, COALA) proposes a numerically robust inversion-free framework for low-rank weighted approximations for this setting. Another idea is to initialize LoRA with a subset of left and right singular vectors of a doubled-rank truncated SVD of the loss function gradient at the starting parameters, proposed by Wang et al. (2024, LoRA-GA). We show direct connection of this method to our Riemannian initialization strategy and propose how to additionally significantly accelerate the computation of SVD using our approach. Attempting to overcome the asymmetry in the initialization of vanilla LoRA fine-tuning process, (Hayou et al., 2024, LoRA+) introduced a scale-free step size selection for LoRA factors.
@@
Riemannian optimization is widely used for algorithms on matrix manifolds and allows for exploiting task geometry or imposing additional constraints. For example, a Riemannian solution for the extreme eigenpairs search problem was described in Absil et al. (2009); Baker (2008), a matrix completion task, which is common in collaborative filtering for recommender systems, via optimization on the fixed-rank manifold (Vandereycken, 2013), a Riemannian approach on the manifold of matrices with orthonormal columns (the Stiefel manifold) was used by Wisdom et al. (2016) for diminishing the problem of vanishing and exploding gradients in recurrent neural networks, etc. The book Trendafilov & Gallo (2021) also presents a comprehensive description of useful manifolds for solutions of the data science problems. For the deeper understanding of applied differential geometry techniques see the books Absil et al. (2009) and Boumal (2023).
@@
The idea of using Riemannian optimization has recently started to emerge for the large language models. For example, the fine-tuning of LLMs with the help of the Stiefel manifold was considered in the work Hu et al. (2024). The authors of (Zhang & Pilanci, 2024) introduced the Riemannian inspired modification of Adam. The authors of Mo et al. (2025, LORO) applied the Riemannian optimization techniques for pretraining LLMs on the fixed-rank manifold. Parametrization that is used in our work can potentially help in this setting as well, by additionally avoiding potential overheads and instabilities, arising due the explicit inversion of Gram matrices.
3 Preliminaries
3.1 Muon optimizer
Muon is an optimizer designed specifically for matrix‑valued parameters in a network’s hidden layers. Empirically, it accelerates training on language and vision workloads while leaving scalar/vector parameters and the input/output layers to a conventional optimizer such as AdamW. At a high level, Muon takes the step that stochastic gradient descent with momentum (SGDM) would make on a weight matrix and orthogonalizes that update before applying it. Orthogonalization acts as a per‑layer, per‑step preconditioner that equalizes singular values of the update, which mitigates the collapse of updates into a few dominant directions (Jordan et al., 2024). More specifically, let $W\in\mathbb{R}^{n\times m}$ be a hidden‑layer weight. With gradient $G_{t}=\nabla_{W}\mathcal{L}(W_{t})$ and momentum $M_{t}=\beta M_{t-1}+G_{t}$, Muon computes
$\widetilde{M}_{t}\;\approx\;\mathrm{Ortho}(M_{t})\quad\text{and}\quad W_{t+1}=W_{t}-\eta\,\widetilde{M}_{t},$
where $\mathrm{Ortho}(\cdot)$ denotes the nearest semi‑orthogonal matrix in Frobenius norm, i.e.,
$\mathrm{Ortho}(G)\;=\;\arg\max_{O}\bigl\{\|O-G\|_{F}:\ O^{\top}O=I\ \text{or}\ OO^{\top}=I\bigr\}.$ (2)
Computing $\mathrm{Ortho}(G)$ exactly amounts to taking the SVD $G=USV^{\top}$ and returning $UV^{\top}$, which is too slow to do at every iteration. Muon instead applies a Newton–Schulz (NS) iteration that—after normalizing $G$ — implements a composition of a fixed low‑degree polynomial in $GG^{\top}$ acting on $G$ and converges to $UV^{\top}$. Leveraging efficient matrix multiplication operations results in a highly performant iteration. We will write $\widetilde{M}_{t}=NS(M_{t})$ for short.
LMO interpretation.
Muon’s step admits a clean linear minimization oracle (LMO) interpretation (Bernstein, 2025). Indeed, consider the operator-norm unit ball $\mathcal{B}_{2}=\{X:\|X\|_{2}\leq 1\}$. The linear minimization oracle (LMO) over $\mathcal{B}_{2}$ at matrix $M_{t}$ given by its SVD $M_{t}=USV^{\top}$ is
$UV^{\top}\in\operatorname*{Arg\,max}_{\|S\|_{2}\leq 1}\ \langle M_{t},S\rangle.$ (3)
Applying Muon to LoRA.
In LoRA, a frozen weight $W_{0}\in\mathbb{R}^{n\times m}$ is adapted via a low‑rank update $W=W_{0}+\alpha\,BA$ with $B\in\mathbb{R}^{n\times r}$, $A\in\mathbb{R}^{r\times m}$ and small $r$. Each trainable factor ($A$ and $B$) is a 2D parameter, so Muon can be applied per factor:
[…]
Note that acting on the two factors separately makes Muon non–reparameterization‑invariant: its per‑factor orthogonalization depends on arbitrary scalings or rotations, skewing the weight‑space step and often letting one factor dominate.
3.2 Riemannian optimization
Let $\mathcal{M}_{r}=\{X\in\mathbb{R}^{m\times n}:\operatorname{rank}(X)=r\}\subseteq\mathbb{R}^{m\times n}$ be a smooth manifold of fixed-rank matrices (Lee, 2003, Example 8.14). Let every point $X$ of $\mathcal{M}_{r}$ be equipped with a tangent plane $\mathcal{T}_{X}\mathcal{M}_{r}$. Thinking geometrically, the tangent plane plays the role of the best local, flat approximation to this curved set: if you “zoom in” at $X$, the manifold looks like a plane. We will now discuss how to numerically parametrize points on a manifold and its tangent plane. Every rank-$r$ matrix $X\in\mathcal{M}_{r}$ can be represented using matrices $A_{L}\in\mathbb{R}^{m\times r},B_{r}\in\mathbb{R}^{n\times r}$ with orthonormal columns and a square matrix $G\in\mathbb{R}^{r\times r}$ as
$X=A_{L}GB_{R}^{\top}.$ (4)
For example, one may think of a thin SVD, in which case $G$ becomes the diagonal matrix of singular values, but other representations are also possible and will be convenient for our purposes. Using (4), any tangent vector $\xi\in\mathcal{T}_{X}\mathcal{M}_{r}$ can be represented in a matrix factorization format:
$\xi=\begin{bmatrix}\dot{A}&A_{L}\end{bmatrix}\begin{bmatrix}B_{R}&\dot{B}\end{bmatrix}^{\top},\quad\dot{A}^{\top}A_{L}=0,$ (5)
@@ 3.2 Riemannian optimization
so that any tangent vector can be identified in terms of the tuple: $(\dot{A},\dot{B})$. Since the factor matrices contain $2r$ columns, we immediately have that $\mathrm{rank}\,\xi\leq 2r$. Another remarkable fact is that the point $X\in\mathcal{M}_{r}$ itself lies in the $\mathcal{T}_{X}\mathcal{M}_{r}$ with $\dot{A}=0,\dot{B}=B$. Given a matrix $Z\in\mathbb{R}^{m\times n}$, its orthogonal projection $P_{\mathcal{T}_{X}\mathcal{M}_{r}}\,Z$ onto the tangent space $\mathcal{T}_{X}\mathcal{M}_{r}$ (with the parameterization given in (5)) can be computed as follows:
$P_{T_{X}\mathcal{M}_{r}}(Z)=A_{L}A_{L}^{\top}Z\;+\;\left(I-A_{L}A_{L}^{\top}\right)ZB_{R}B_{R}^{\top}.$ (6)
and, hence, can be represented in the form of (5) with $\dot{A}=(I-A_{L}A_{L}^{\top})ZB_{R}$, and $\dot{B}=Z^{\top}A_{L}$.
@@
Let $F\colon\mathbb{R}^{m\times n}\to\mathbb{R}$ be a differentiable function with the Euclidean gradient $\nabla F\in\mathbb{R}^{m\times n}$. Within the Riemannian optimization framework, we solve the following task optimization problem: